cs Scanned with CamScanner

2. A helical compression spring is to be designed as shown in Fig Q2. The spring is required to exert a force of Po-102 N when compressed to a length of lo=44 mm. When its length is 1=76 mm, it must exert a force of P=22 N. The spring will be cycled rapidly, with severe service required. Use ASTM A401 steel wire. The factor of safety, η, need to be larger than 1.5. Starting with a spring index of 9, and checking for stability is required. a) Compute the spring rate k and the free length Ir of the spring. (6 pts) b) Determine the wire diameter dw, of the spring to satisfy the conditions given, assuming a peened spring. Use the Goodman failure locus for the Zimmerli data. (11 pts) c) Determine the solid length s, the force Ps and stress corresponding to the solid length, the number of active coils, OD of the coil and the factor of safety guarding against static failure. (18 pts)
cs Scanned with CamScanner
Answer #1

L>1,5 Pnnj Ca Calciatz 3 P*ropo Sove ,2 equaion 22 - Klf 44k 80 32K 22 - 215 Cuf- 8.8 - Lf-44 = 44-8.8 35.2 mm Spring cota sFOS HC-4 ven c 36-I 36-4 Fos -fado of safety K 35093 32 fom-m으 ptoperty data sheet of ASTM A401 steel FOS FOS 29 こ 0.5チチSut 43 siu diametan s 2.9 mm CDetemonofNumben ofaie ih s cocua usina ialalicn Gid 8PD where ổ: spring 2.S Modda o81310 x219 SubstiSpinq a 52, 2 mm ce s - ,5 x 52,9 hu covTesponding to soid lanth aI 21 = 26-1,十 2